Solve by completing the square:
2x^2-8x+1=0

Answer by stanbon(75887)

Solve by completing the square:
2x^2-8x+1=0
-----
x^2 - 4x + ? = -1/2 + ?
-----
x^2 -4x + 4 = (-1/2) + 4
---
(x-2)^2 = 7/2
---
x-2 = sqrt(14)/2 or x-2 = -sqrt(14)/2
---
x = 2 + (1/2)sqrt(14) or x = 2 - (1/2(sqrt(14)
=======================
Cheers,
Stan H.
================